# XRF V2: A Dataset for Action Summarization with Wi-Fi Signals, and IMUs in Phones, Watches, Earbuds, and Glasses

### v1

@@ Abstract.
Human Action Recognition (HAR) plays a crucial role in applications such as health monitoring, smart home automation, and human-computer interaction. While HAR has been extensively studied, action summarization, which involves identifying and summarizing continuous actions, remains an emerging task. This paper introduces the novel XRF V2 dataset, designed for indoor daily activity Temporal Action Localization (TAL) and action summarization. XRF V2 integrates multimodal data from Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings, offering a diverse collection of indoor activities from 16 volunteers across three distinct environments. To tackle TAL and action summarization, we propose the XRFMamba neural network, which excels at capturing long-term dependencies in untrimmed sensory sequences and outperforms state-of-the-art methods, such as ActionFormer and WiFiTAD. We envision XRF V2 as a valuable resource for advancing research in human action localization, action forecasting, pose estimation, multimodal foundation models pre-training, synthetic data generation, and more.
@@
copyright: acmlicensedjournalyear: 2025publicationmonth: 1doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Human Indoor Action Understanding (HIAU) is pivotal in enabling intelligent environments, with wide-ranging applications in health monitoring, smart home automation, and human-computer interaction. Currently, human action recognition (HAR) is the most extensively studied aspect in HIAU. HAR aims to identify and categorize discrete human actions by leveraging data from diverse modal sensors, such as smartwatches (Laput and Harrison, 2019; Wang et al., 2024b), smartphones (Sun et al., 2024; DeVrio et al., 2023), earbuds (Prakash et al., 2020; Zhao et al., 2024), smart glasses (Lee and Hui, 2018; Xie et al., 2021), and Wi-Fi signals (Ma et al., 2018; Wang et al., 2016a, 2017a). With accurate HAR, systems can anticipate user needs, improve safety, and provide context-aware assistance, fostering a seamless interaction between individuals and their surroundings.
In this paper, we take a step forward to propose a new HIAU task: action summarization. As shown in Fig. 1, users perform continuous actions indoors, which are identified by recognizing each action’s category, start time, and end time to form a sequence of tuples. These tuples can be processed using large language models (LLMs) to generate action summaries, such as insights into physical activity levels and behavioral patterns. With targeted prompts, the system can also detect anomalies like falls or missed medication. We envision several applications for action summarization: (1) Health Monitoring: providing detailed activity logs to support physical rehabilitation, track fitness goals, or detect health issues such as sedentary behavior or irregular movement patterns. (2) Elderly Care: detecting critical events like falls, identifying missed medication times, and offering behavior insights to caregivers for better elderly support. (3) Smart Home Automation: triggering adaptive responses, such as adjusting room temperature based on activity levels or controlling lighting when specific actions are performed. (4) Personal Assistants: Summarizing daily routines to help users with task planning, reminders, and prioritizing activities, thereby enhancing productivity.
@@
To address these limitations, we present XRF V2, a novel dataset designed specifically for summarizing indoor daily activities. XRF V2 integrates data from multiple sensing modalities, including Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings, as illustrated in Fig. 1. This dataset was collected with the participation of 16 volunteers, who performed diverse and continuous action sequences across three distinct indoor environments: the dining room, study room, and bedroom, aiming to support a wide range of applications in action summarization. In total, XRF V2 consists of approximately 16 hours and 16 minutes of multimodal data recordings, comprising 853 annotated action sequences that encompass diverse daily routines, including sleeping, cutting fruit, and taking medicine. Compared to WiFiTAD (Liu et al., 2025), XRF V2 provides a more diverse, realistic, and comprehensive benchmark for the action summarization task.
@@
We further propose the XRFMamba neural network to evaluate the XRF V2 dataset and conduct accurate Temporal Action Localization (TAL) for action summarization. XRFMamba is based on the latest advanced achievements in Mamba, which excels at capturing long-term dependencies in temporal sequences. We choose XRFMamba because, unlike action recognition, TAL involves longer input sequences that contain multiple actions, and Mamba has already demonstrated strong capabilities in handling long-term dependencies in fields like NLP (Gu and Dao, 2023) and CV (Zhu et al., 2024). For example, ActionMamba (Chen et al., 2024) has been successfully applied for video-based TAL tasks. In this paper, we are the first to apply the Mamba network to handle Wi-Fi and IMU data for TAL tasks. Compared with existing methods such as ActionFormer (Zhang et al., 2022), TriDet (Shi et al., 2023), ActionMamba (Chen et al., 2024), TemporalMaxer (Tang et al., 2023), UWiFiAction (Wang et al., 2023), and WiFiTAD (Liu et al., 2025), the proposed XRFMamba achieves superior results, e.g., outperforming the latest WiFiTAD by a large margin of 5.49 points in mAP, while using 35% fewer parameters.
@@
(2) New Task, Method, and Evaluation Metric: We introduce the novel task of action summarization, a new human indoor action understanding task. We also propose a new evaluation metric, Response Meaning Consistency (RMC), to assess the performance of action summarization. We further design a novel method, XRFMamba, to effectively tackle this task. Compared to existing state-of-the-art methods, XRFMamba demonstrates superior performance in action summarization.
@@ 2.1. Action Understanding with Wi-Fi Signals
Action understanding with Wi-Fi signals has become a key technology in human sensing. The RSSI (Received Signal Strength Indicator) in Wi-Fi signals is commonly used for indoor localization (Bahl and Padmanabhan, 2000; Wu et al., 2012b). However, its accuracy is relatively low and is susceptible to multipath effects and temporal dynamics (Yang et al., 2013; Wu et al., 2011). As a result, later works have adopted more information-rich Channel State Information (CSI) to achieve more precise localization (Yang et al., 2013), such as FILA (Wu et al., 2012a), SpotFi (Kotaru et al., 2015), Splicer (Xie et al., 2015), and Dynamic-MUSIC (Li et al., 2016a). In addition to high-precision indoor localization, Wi-Fi CSI signals have been widely applied in various tasks in fields such as smart homes, health monitoring, and human-computer interaction. These include hand gesture recognition (Pu et al., 2013; Ma et al., 2018), finger gesture recognition (Li et al., 2016b; Tan and Yang, 2016; Ali et al., 2015), activity recognition (Wang et al., 2015, 2016a, 2014), vital sign monitoring (Wang et al., 2016b, 2017b; Yu et al., 2021), pose estimation (Jiang et al., 2020; Wang et al., 2019b; Yan et al., 2024; Wang et al., 2019a), mesh reconstruction (Wang et al., 2022b, c), crowd counting (Xi et al., 2014), and human imaging (Xu et al., 2024; Wang et al., 2019b). Since Wi-Fi-based human sensing is highly sensitive to environmental changes, several studies have concentrated on enhancing its cross-environment capabilities. For example, EI (Jiang et al., 2018) and Person-in-WiFi (Wang et al., 2019b) use domain adaptation techniques to align the features of the source and target environments for activity recognition and pose estimation, respectively. CrossSense (Zhang et al., 2018) proposes using a mixture of experts approach to train multiple tasks and select the optimal expert for the target environment, maintaining cross-environment adaptability. Widar 3.0 (Zheng et al., 2019), on the other hand, proposes environment-invariant representation, i.e., body velocity profiles, to achieve cross-environment, cross-person, and cross-method hand gesture recognition.
@@
Apart from the tasks mentioned above, some researchers have also focused on understanding continuous action using Wi-Fi. Smokey (Zheng et al., 2016) recognizes smoking actions by detecting periodic patterns, such as holding the cigarette, bringing it to the mouth, and inhaling the smoke. WiLife (Li et al., 2024a) uses Wi-Fi for continuous in-home monitoring of elderly individuals living alone, ensuring their safety and well-being. Wang et al. (Wang et al., 2023) employ U-shaped deep neural networks for the temporal localization and segmentation of daily activities and human-human interactions. WiFiTAD (Liu et al., 2025) introduces a frequency-aware learning approach combined with a dual-pyramid network to promote the accuracy of temporal action localization. IMar (He and Yang, 2022) employs tensor decomposition to extract individual Wi-Fi streams for each person, enabling separate recognition of continuous actions. In comparison to these works, we push the boundaries further by introducing a new task: action summarization. We envision this task supporting applications in remote healthcare and monitoring, personal assistants, smart home automation, and more.
@@ 2.2. Action Understanding with IMUs
Inertial Measurement Units (IMUs), such as accelerometers and gyroscopes, can measure motion acceleration and angular velocity. They are widely embedded in personal consumer electronics, including smartphones, smartwatches, smart glasses, and earbuds. Researchers have leveraged IMUs in these devices to develop numerous applications in areas such as healthcare (Alam, 2020; Jung et al., 2022; DeVrio et al., 2023; Shang et al., 2024; Wang et al., 2024b; De Brabandere et al., 2020; Wang et al., 2022d; Meng et al., 2022), sports (Tian et al., 2021; Wolff et al., 2018), human-computer interaction (Laput and Harrison, 2019; DeVrio et al., 2023; Sun et al., 2024; DeVrio et al., 2023; Wahl et al., 2015; Scholl et al., 2015; Verma et al., 2021; Zhao et al., 2024), and industrial applications (Azadi et al., 2019; Sopidis et al., 2022). Jung et al. (Jung et al., 2022) utilize smartwatches and continuous wavelet transform (CWT) technology to identify tremor episodes and amplitude of Parkinson’s disease patients during drawing tasks. Azadi et al. (Azadi et al., 2019) investigate the feasibility of recognizing screwing actions in manufacturing using smartwatches and unsupervised learning techniques. De et al. (De Brabandere et al., 2020) place a mobile phone on a patient’s left hip to estimate hip and knee joint loading, providing valuable insights for doctors in diagnosing the patient’s condition. Anguita et al. (Anguita et al., 2013) leverage a smartphone worn at the waist to identify six daily activities: standing, sitting, lying down, walking, walking downstairs, and walking upstairs. WISEGlass (Wahl et al., 2015) captures the wearer’s head movements—up, right, left, and down—using IMUs embedded in the legs of smart glasses, enabling the control of the Pac-Man game accordingly. Wetlab (Scholl et al., 2015) integrates Google Glass with a wrist-worn IMU sensor, allowing for the capture and review of DNA extraction experiments conducted in wet laboratory environments. ExpressEar (Verma et al., 2021) uses earbuds to detect fine facial muscle movements, exploring the correlation between movements and facial expressions. This enables hands-free control through intentional facial expressions or monitoring involuntary ones to infer the user’s emotional state. ExpressEar (Verma et al., 2021) utilizes a Nokia eSense earbuds to sense fine facial muscle movements, and mine the correlation between the movements and facial expressions for hands-free control using facial expressions or to monitor involuntary facial expressions to infer the user’s emotional state. Ui-Ear (Zhao et al., 2024) utilizes earbuds to detect on-face gestures, including tapping, slide up, slide down, slide left, slide right, clockwise circle, and counterclockwise circle, enhancing the user-earbuds interaction experience.
@@ 2.3. Continuous Action Dataset and Benchmark
Temporal action localization (TAL) and temporal action detection (TAD) are fundamental tasks in the computer vision community and have been extensively studied (Shou et al., 2016). Searchers have introduced numerous publicly accessible datasets to evaluate TAL and TAD methods. Breakfast (Kuehne et al., 2014) and YouCook2 (Zhou et al., 2018) are two representative large-scale daily cooking datasets captured from third-person viewpoint cameras. The Breakfast includes 10 operations related to breakfast preparation, performed by 52 different people from 18 different kitchens. YouCook2 contains 2000 untrimmed long videos from 89 cooking recipes, with each step annotated with temporal boundaries and described in sentences. Beyond these cooking datasets, some other datasets such as ActivityNet (Caba Heilbron et al., 2015), THUMOS14 (Idrees et al., 2017), HACS Segments (Zhao et al., 2019), FineAction (Liu et al., 2022), are released for TAL involving daily activities, sports, and household tasks. Among these, FineAction consists of 103K temporal instances of 106 action categories, annotated in 17K untrimmed videos, covering household activities, personal care, socializing, and sports. While cameras are widely used for TAL tasks, they face challenges such as fixed positions and viewing angles once deployed, as well as privacy concerns in home environments. Thus, several studies have explored smartwatches, smartphones, and Wi-Fi for TAL. UWash (Wang et al., 2024b) is a dedicated system for segmenting handwashing gestures using smartwatch IMU data. Its dataset includes 550 handwashing recordings from 71 participants. Ego-ADL (Sun et al., 2024) leverages a smartphone to continuously capture users’ audio, Wi-Fi, and motion sensor data in daily life. The released dataset comprises 141 annotated sequences. However, the reliance on a single smartphone placed in the user’s pocket introduces limitations in capturing fine-grained hand and head movements. WiFiTAD (Liu et al., 2025) uses a setup with one transmitter and one receiver Wi-Fi device to collect data in a single indoor environment. The dataset features recordings from three participants performing seven daily activities—walking, running, jumping, waving, falling, sitting, and standing. It includes 553 untrimmed sequences, with an average of 3.79 action instances per sequence.
@@
XRF V2 3 30 15 853 (16h 16m 8s) Wi-Fi, 5 IMUs, RGB+D+IR /

### v2

@@ Abstract.
Human Action Recognition (HAR) plays a crucial role in applications such as health monitoring, smart home automation, and human-computer interaction. While HAR has been extensively studied, action summarization using Wi-Fi and IMU signals in smart-home environments , which involves identifying and summarizing continuous actions, remains an emerging task. This paper introduces the novel XRF V2 dataset, designed for indoor daily activity Temporal Action Localization (TAL) and action summarization. XRF V2 integrates multimodal data from Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings, offering a diverse collection of indoor activities from 16 volunteers across three distinct environments. To tackle TAL and action summarization, we propose the XRFMamba neural network, which excels at capturing long-term dependencies in untrimmed sensory sequences and achieves the best performance with an average mAP of 78.74, outperforming the recent WiFiTAD by 5.49 points in mAP@avg while using 35% fewer parameters. In action summarization, we introduce a new metric, Response Meaning Consistency (RMC), to evaluate action summarization performance. And it achieves an average Response Meaning Consistency (mRMC) of 0.802. We envision XRF V2 as a valuable resource for advancing research in human action localization, action forecasting, pose estimation, multimodal foundation models pre-training, synthetic data generation, and more. The data and code are available at https://github.com/aiotgroup/XRFV2.
@@
* Fei Wang is the corresponding author who initiated and led the project. Bo Lan and Pei Li contributed equally.
copyright: acmlicensedjournal: IMWUTjournalyear: 2025publicationmonth: 9journalvolume: 0journalnumber: 0doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Figure 1. XRF V2 includes multimodal data from Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings. Volunteers perform continuous actions indoors, and Temporal Action Localization (TAL) techniques are used to identify the start and end times of each action within untrimmed sequences, which are then combined into tuples. These action sequence tuples, in conjunction with task-oriented prompts, are input into LLMs, thereby enabling the LLMs to function as intelligent agents, supporting various applications, such as remote healthcare monitoring, personal assistant, smart home automation, and health and wellness analysis.
Human Indoor Action Understanding (HIAU) is pivotal in enabling intelligent environments, with wide-ranging applications in health monitoring, smart home automation, and human-computer interaction. Currently, human action recognition (HAR) is the most extensively studied aspect in HIAU. HAR aims to identify and categorize discrete human actions by leveraging data from diverse modal sensors, such as smartwatches (Laput and Harrison, 2019; Wang et al., 2025), smartphones (Sun et al., 2024; DeVrio et al., 2023), earbuds (Prakash et al., 2020; Zhao et al., 2024), smart glasses (Lee and Hui, 2018; Xie et al., 2021), and Wi-Fi signals (Ma et al., 2018; Wang et al., 2016a, 2017a). With accurate HAR, systems can anticipate user needs, improve safety, and provide context-aware assistance, fostering a seamless interaction between individuals and their surroundings.
In this paper, we take a step forward to propose a new HIAU task: action summarization using Wi-Fi and IMU signals in smart-home environments. As shown in Fig. 1, users perform continuous actions indoors, which are identified by recognizing each action’s category, start time, and end time to form a sequence of tuples. These tuples can be processed using large language models (LLMs) to generate action summaries, such as insights into physical activity levels and behavioral patterns. With targeted prompts, the system can also detect anomalies like falls or missed medication. We envision several applications for action summarization: (1) Health Monitoring: providing detailed activity logs to support physical rehabilitation, track fitness goals, or detect health issues such as sedentary behavior or irregular movement patterns. (2) Elderly Care: detecting critical events like falls, identifying missed medication times, and offering behavior insights to caregivers for better elderly support. (3) Smart Home Automation: triggering adaptive responses, such as adjusting room temperature based on activity levels or controlling lighting when specific actions are performed. (4) Personal Assistants: Summarizing daily routines to help users with task planning, reminders, and prioritizing activities, thereby enhancing productivity.
@@
To address these limitations, we present XRF V2, a novel dataset designed specifically for summarizing indoor daily activities. XRF V2 integrates data from multiple sensing modalities, including Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings, as illustrated in Fig. 1. XRF V2 combines Wi-Fi and IMU sensors to capture human motions, providing a more comprehensive understanding of actions. In real-world scenarios, sensor devices may experience malfunctions. For example, if a smartphone’s IMU malfunctions or the device is not worn as intended (e.g., left on a table), the IMU data may no longer accurately represent the user’s actions. Similarly, a malfunctioning Wi-Fi router can disrupt Wi-Fi signals. XRF V2 is designed to handle such situations by adapting to different combinations of sensor inputs for Temporal Action Localization (TAL) and action summarization. This flexibility ensures the system remains functional even when some sensors are not operating optimally. Moreover, Wi-Fi and IMU sensors are commonly found in commercial products, making it easy to integrate TAL and action summarization systems into existing home and personal devices without the need for costly specialized equipment. This dataset was collected with the participation of 16 volunteers, who performed diverse and continuous action sequences across three distinct indoor environments: the dining room, study room, and bedroom, aiming to support a wide range of applications in action summarization using Wi-Fi and IMU signals in smart-home environments. In total, XRF V2 consists of approximately 16 hours and 16 minutes of multimodal data recordings, comprising 853 annotated action sequences that encompass diverse daily routines, including sleeping, cutting fruit, and taking medicine. Compared to WiFiTAD (Liu et al., 2025), XRF V2 provides a more diverse, realistic, and comprehensive benchmark for the action summarization task.
@@
We further propose the XRFMamba neural network to evaluate the XRF V2 dataset and conduct accurate Temporal Action Localization (TAL) for action summarization. XRFMamba is based on the latest advanced achievements in Mamba, which excels at capturing long-term dependencies in temporal sequences. We choose XRFMamba because, unlike action recognition, TAL involves longer input sequences that contain multiple actions, and Mamba has already demonstrated strong capabilities in handling long-term dependencies in fields like NLP (Gu and Dao, 2023) and CV (Zhu et al., 2024). For example, ActionMamba (Chen et al., 2024) has been successfully applied for video-based TAL tasks. In this paper, we are the first to apply the Mamba network to handle Wi-Fi and IMU data for TAL tasks. Compared with existing methods such as ActionFormer (Zhang et al., 2022), TriDet (Shi et al., 2023), ActionMamba (Chen et al., 2024), TemporalMaxer (Tang et al., 2023), UWiFiAction (Wang et al., 2023), and WiFiTAD (Liu et al., 2025), the proposed XRFMamba excels at capturing long-term dependencies in untrimmed sensory sequences and achieves the best performance with an average mAP of 78.74, outperforming the recent WiFiTAD by 5.49 points in mAP@avg while using 35% fewer parameters. In action summarization, we introduce a new metric, Response Meaning Consistency (RMC), to evaluate action summarization. And it achieves an average Response Meaning Consistency(mRMC) of 0.802.
Moreover, we conducted a leave-one-person-out evaluation, where XRFMamba achieved a mAP of 72.84 for new users, demonstrating strong generalization capabilities. We also performed device combination experiments to assess the model’s flexibility under different device configurations. Notably, even when using only a single right-hand-worn smartwatch, XRFMamba still achieved a mAP of 65.81, indicating that users can flexibly balance performance and device availability. Additionally, we tested XRFMamba’s real-world deployment potential in a real apartment, where the system achieved an mAP of around 70, further validating its practical applicability.
@@
(2) New Task, Method, and Evaluation Metric: We introduce the novel task of action summarization using Wi-Fi and IMU signals in smart-home environments, a new human indoor action understanding task. We also propose a new evaluation metric, Response Meaning Consistency (RMC), to assess the performance of action summarization. We further design a novel method, XRFMamba, to effectively tackle this task. Compared to existing state-of-the-art methods, XRFMamba demonstrates superior performance in action summarization.
@@ 2.1. Action Understanding with Wi-Fi Signals
Action understanding with Wi-Fi signals has become a key technology in human sensing. The RSSI (Received Signal Strength Indicator) in Wi-Fi signals is commonly used for indoor localization (Bahl and Padmanabhan, 2000; Wu et al., 2012b). However, its accuracy is relatively low and is susceptible to multipath effects and temporal dynamics (Yang et al., 2013; Wu et al., 2011). As a result, later works have adopted more information-rich Channel State Information (CSI) to achieve more precise localization (Yang et al., 2013), such as FILA (Wu et al., 2012a), SpotFi (Kotaru et al., 2015), Splicer (Xie et al., 2015), and Dynamic-MUSIC (Li et al., 2016a). In addition to high-precision indoor localization, Wi-Fi CSI signals have been widely applied in various tasks in fields such as smart homes, health monitoring, and human-computer interaction. These include hand gesture recognition (Pu et al., 2013; Ma et al., 2018), finger gesture recognition (Li et al., 2016b; Tan and Yang, 2016; Ali et al., 2015), activity recognition (Wang et al., 2015, 2016a, 2014), vital sign monitoring (Wang et al., 2016b, 2017b; Yu et al., 2021), pose estimation (Jiang et al., 2020; Wang et al., 2019c; Yan et al., 2024; Wang et al., 2019b), mesh reconstruction (Wang et al., 2022b, c), crowd counting (Xi et al., 2014), and human imaging (Xu et al., 2024; Wang et al., 2019c). Since Wi-Fi-based human sensing is highly sensitive to environmental changes, several studies have concentrated on enhancing its cross-environment capabilities. For example, EI (Jiang et al., 2018) and Person-in-WiFi (Wang et al., 2019c) use domain adaptation techniques to align the features of the source and target environments for activity recognition and pose estimation, respectively. CrossSense (Zhang et al., 2018) proposes using a mixture of experts approach to train multiple tasks and select the optimal expert for the target environment, maintaining cross-environment adaptability. Widar 3.0 (Zheng et al., 2019), on the other hand, proposes environment-invariant representation, i.e., body velocity profiles, to achieve cross-environment, cross-person, and cross-method hand gesture recognition.
@@
Apart from the tasks mentioned above, some researchers have also focused on understanding continuous action using Wi-Fi. Smokey (Zheng et al., 2016) recognizes smoking actions by detecting periodic patterns, such as holding the cigarette, bringing it to the mouth, and inhaling the smoke. WiLife (Li et al., 2024a) uses Wi-Fi for continuous in-home monitoring of elderly individuals living alone, ensuring their safety and well-being. Wang et al. (Wang et al., 2023) employ U-shaped deep neural networks for the temporal localization and segmentation of daily activities and human-human interactions. WiFiTAD (Liu et al., 2025) introduces a frequency-aware learning approach combined with a dual-pyramid network to promote the accuracy of temporal action localization. IMar (He and Yang, 2022) employs tensor decomposition to extract individual Wi-Fi streams for each person, enabling separate recognition of continuous actions. In comparison to these works, we push the boundaries further by introducing a new task: action summarization using Wi-Fi and IMU signals in smart-home environments. We envision this task supporting applications in remote healthcare and monitoring, personal assistants, smart home automation, and more.
@@ 2.2. Action Understanding with IMUs
Inertial Measurement Units (IMUs), such as accelerometers and gyroscopes, can measure motion acceleration and angular velocity. They are widely embedded in personal consumer electronics, including smartphones, smartwatches, smart glasses, and earbuds. Researchers have leveraged IMUs in these devices to develop numerous applications in areas such as healthcare  (Alam, 2020; Jung et al., 2022; DeVrio et al., 2023; Shang et al., 2024; Wang et al., 2025; De Brabandere et al., 2020; Wang et al., 2022d; Meng et al., 2022), sports (Tian et al., 2021; Wolff et al., 2018), human-computer interaction (Laput and Harrison, 2019; DeVrio et al., 2023; Sun et al., 2024; DeVrio et al., 2023; Wahl et al., 2015; Scholl et al., 2015; Verma et al., 2021; Zhao et al., 2024), and industrial applications  (Azadi et al., 2019; Sopidis et al., 2022). Jung et al. (Jung et al., 2022) utilize smartwatches and continuous wavelet transform (CWT) technology to identify tremor episodes and amplitude of Parkinson’s disease patients during drawing tasks. Azadi et al. (Azadi et al., 2019) investigate the feasibility of recognizing screwing actions in manufacturing using smartwatches and unsupervised learning techniques. De et al. (De Brabandere et al., 2020) place a mobile phone on a patient’s left hip to estimate hip and knee joint loading, providing valuable insights for doctors in diagnosing the patient’s condition. Anguita et al. (Anguita et al., 2013) leverage a smartphone worn at the waist to identify six daily activities: standing, sitting, lying down, walking, walking downstairs, and walking upstairs. WISEGlass (Wahl et al., 2015) captures the wearer’s head movements—up, right, left, and down—using IMUs embedded in the legs of smart glasses, enabling the control of the Pac-Man game accordingly. Wetlab (Scholl et al., 2015) integrates Google Glass with a wrist-worn IMU sensor, allowing for the capture and review of DNA extraction experiments conducted in wet laboratory environments. ExpressEar (Verma et al., 2021) utilizes a Nokia eSense earbuds to sense fine facial muscle movements, and mine the correlation between the movements and facial expressions for hands-free control using facial expressions or to monitor involuntary facial expressions to infer the user’s emotional state. Ui-Ear (Zhao et al., 2024) utilizes earbuds to detect on-face gestures, including tapping, slide up, slide down, slide left, slide right, clockwise circle, and counterclockwise circle, enhancing the user-earbuds interaction experience.
@@ 2.3. Continuous Action Dataset and Benchmark
XRF V2 3 30 15 853 (16h 16m 8s) Wi-Fi, 5 IMUs, RGB+D+IR IMWUT 2025
Temporal action localization (TAL) and temporal action detection (TAD) are fundamental tasks in the computer vision community and have been extensively studied (Shou et al., 2016). Searchers have introduced numerous publicly accessible datasets to evaluate TAL and TAD methods. Breakfast (Kuehne et al., 2014) and YouCook2 (Zhou et al., 2018) are two representative large-scale daily cooking datasets captured from third-person viewpoint cameras. Beyond these cooking datasets, some other datasets such as ActivityNet (Caba Heilbron et al., 2015), THUMOS14 (Idrees et al., 2017), HACS Segments (Zhao et al., 2019), FineAction (Liu et al., 2022), are released for TAL involving daily activities, sports, and household tasks. UWash (Wang et al., 2025) is a dedicated system for segmenting handwashing gestures using smartwatch IMU data. Ego-ADL (Sun et al., 2024) leverages a smartphone to continuously capture users’ audio, Wi-Fi, and motion sensor data in daily life. WiFiTAD (Liu et al., 2025) uses a setup with one transmitter and one receiver Wi-Fi device to collect data in a single indoor environment.